# An Adaptive Placement and Parallelism Framework for Accelerating RLHF Training

### 19 Dec 2023 v1

@@ V Our Framework
V-A Model Placement Strategies
Figure 4: The timeline of Flattening strategy vs. Interleaving strategy. Two homogeneous GPU devices are utilized for demonstration. Under the Flattening strategy, both the Reference model and Reward model are deployed on both devices, whereas under the Interleaving strategy, the Reference model and Reward model are allocated to separate devices.
V-A1 Interleaving Strategy
@@
The main idea of the Interleaving strategy is to partially parallelize the pipeline by allocating the independent models on exclusive group of GPU devices, as well as shrinking the number of computing devices required for each model. We observe that either Reward&Reference Inference (and Actor&Critic training models) could operate independently, thereby it is possible to eliminate part of the serial execution. For example, the Reference model and Reward model both take the generated responses from the Actor model as input and output the intermediate results into the experience buffer without synchronization. Also, the Actor and Critic do not need synchronization duration of the training phase. This design helps the parallelization of independent models in time-cost RLHF training. This lowers the memory redundancy and shrinks the scale of participating devices used by either the Reference or Reward model rather than using all devices in the Flattening strategy.
@@
We compare the distributed architecture and computation procedure between the Flattening strategy and our Interleaving strategy, using a small number of devices as illustrative examples. Our strategy could be easily extended to large-scale computing devices. In terms of Model Placement, the Flattening strategy treats the four models as a unit and deploys them on all devices, as depicted in Figure  3(a). By contrast, our Interleaving strategy distinctly assigns the Reference model and Reward model to two distinct device groups (GPU 0-3 and 4-7) [^1] Our interleaving strategy could further place the Actor and Critic model on exclusive devices. We only illustrate the interleaving between Reward and Reference for simplicity., as illustrated in Figure  3(b). Our Interleaving placement strategy reduces the number of devices used by either the Reference model or the Reward model. It’s important to note that responses generated from the Actor are located on GPUs where the Reference model is not placed since the Actor model is deployed on all GPU devices, as shown in the timeline in Figure  4(b). Therefore, it is necessary to perform two extra communication stages: an AllGather primitive to gather the generated results from all devices before forwarding, and then perform a Scatter primitive to distribute the results back to all devices. However, these two additional communication stages can be considered negligible due to overlapping computation and communication. We conclude that the Interleaving strategy shortens the time-cost generation duration from two aspects.
@@
Firstly, in terms of memory consumption, our Interleaving strategy reduces the number of participating devices for either the Reference or Reward model by half, effectively minimizing memory redundancy, especially in ZeRO-0 parallelism. This allows our method to leverage the saved memory to increase the total batch sizes, resulting in higher overall throughput. For instance, when employing the Flattening strategy under ZeRO-0, all devices hold replicas of the model states, while only half of devices need to hold the replicas under our Interleaving strategy. As depicted in Figure  4(b), our method can accommodate a larger batch size compared to the Flattening strategy, as shown in Figure  4(a).
Figure 5: The timeline of Separation strategy on homogeneous (a) and heterogeneous GPU devices (b). In the homogeneous Separation strategy, the training and inference models are placed on GPU 1&2 and GPU 3&4 respectively. In the heterogeneous Separation strategy, these two stages are allocated to dedicated devices exclusively for either training or inference purposes.
@@
Secondly, from a communication efficiency perspective, having few devices leads to reduced communication overhead in both the generation and training stages. The communication cost increases directly with the number of participating devices, particularly for inter-node communication. It’s important to note that while increased levels of ZeRO parallelism reduce memory redundancy in model states, it comes at the cost of increased communication[27, 38]. For instance, ZeRO-3 Parallelism allows for accommodating larger models in less memory. However, during both the generation and training stages, an AllGather operation is required to obtain all parameters of the current layer before usage. As the scale of computing nodes increases, the problem is exacerbated, especially in GPU clusters with low-bandwidth inter-node communication, which is the bottleneck in current distributed LLM training. However, the Interleaving strategy can further leverage high-speed intra-node bandwidth and reduce inter-node communication by colocating the Reward model or Reference model on a single machine with usually 8 GPU devices, particularly in small to medium training scenarios. This approach significantly decreases communication overhead, especially inter-node communication, and improves overall efficiency.
Furthermore, we eliminate the Forward step of the Actor model by aggregating it into the Actor generation step, as indicated by the dashed box in Figure  4(a). This aggregation in computation steps also brings certain efficiency gains.
V-A2 Separation Strategy
@@
To shorten the generation duration, our Separation strategy involves separating the training stage and the generation stage of the RLHF pipeline. We recognize that the training and inference stages are inherently different. This division does not only allow us to run the training and generation models simultaneously but also leverage various powerful inference optimization techniques and idle heterogeneous devices in clusters. It could largely accelerate the generation stage and re-balance the workloads between these two stages in RLHF pipeline.
@@
This strategy significantly speeds up the entire RLHF training, although it introduces certain memory redundancy and communication overhead. Specifically, in our strategy, the Actor and Critic models are classified into Prediction(Pred) Actor/Critic models and Training Actor/Critic models by duplicating, which are exclusively used for training or inference. Figure  4(c) illustrates the placement of the training Actor and Critic models and other prediction models, including the Pred Actor and Critic models, on separate devices. Optionally, we could further place the training Actor and Critic models on exclusive devices in an interleaving manner to further decrease the memory redundancy and communication costs. However, there is a need for communication for intermediate results and parameter synchronization between the training Actor/Critic models and the prediction models. Firstly, as shown in Figure  5(a), extra point-to-point communications are required before or after the Forward stage of each prediction model. Additionally, the training Actor and Critic models need to pull the intermediate results from the inference devices. These extra communication stages introduce some idle time per device, known as bubble overhead. However, these bubbles can be minimized by overlapping computation and communication stages through the use of pipeline parallelism. As illustrated in Figure  5(a), the bubbles are nearly hidden through micro-batching in a pipeline parallelism manner, especially with a higher number of generation times. Secondly, it is necessary to synchronize the trained model parameters between the training Actor/Critic models and the inference models after each round of PPO training. The synchronization overhead is inevitable but acceptable.
Although this approach introduces some complexity and overhead, it provides significant benefits from three different perspectives:
Firstly, from the perspective of parallelism, our Separation strategy decouples the training and generation stages for the Actor/Critic models, as discussed in Section IV. This allows us to apply suitable parallelism strategies to either training or the Pred Actor/Critic model. For example, the time-consuming Pred Actor model can be accelerated by utilizing specialized parallelism strategies that are well-suited for inference workloads, such as intra-node tensor parallelism. To be more specific, the Pred Actor model can be replicated on different machines, and the parameters can be sharded horizontally or vertically across the devices within the same machine. This optimization fully takes advantage of the high-bandwidth intra-node communication and minimizes inter-node communication.
Secondly, the separation between the training and generation stages allows the APP framework to easily leverage the rapidly evolving inference optimization techniques by integrating different open-source libraries, such as quantization[7, 41] or speculative sampling[4]. In contrast, under the Flattening strategy, it is challenging to fully take advantage of these optimization techniques due to the usage of the same model for both training and prediction stages. Also, the frequent swapping between training and prediction modes of the same model incurs extra costs.
Lastly, this strategy allows the APP framework to scale out additional computing resources, including heterogeneous devices, for the time-consuming generation stage. This scalability helps rebalance the throughput between the training and generation stages, resulting in accelerated overall training efficiency. With this approach, high-end GPU devices like A100 can be used for training models, while more cost-effective GPU devices such as NVIDIA A10 or even 3090 can be utilized for inference models. This flexibility in either quantity or type of devices offers significant adaptability to diverse training scenarios.
V-B APP Execution Engine
The APP Execution Engine is designed to provide a user-friendly interface for modeling experts to easily build their own RLHF training pipeline. It is compatible with various optimization techniques and supports different model placement strategies to meet the requirements of various scale training scenarios. Our placement strategies may involve 3 to 6 models, including Pred Actor, Pred Critic, Reference, Reward, Train-Actor, and Train-Critic, along with heterogeneous devices based on the chosen placement strategy as well as AC share or AC Non-Share structure.
@@ V-B APP Execution Engine
Model Placement Abstraction. The APP Execution Engine encapsulates each model involved in RLHF pipeline considering both AC share and AC Non-Share structures. Our modular design allows for the encapsulation of each model and corresponding distributed techniques. These techniques include the model placement strategy, singular model parallelism (such as ZeRO in DeepSpeed and Tensor Parallelism in Megatron), as well as various optimization techniques in other libraries. While our placement strategies offer general guidelines on how to distribute these models, there remains a gap in determining the optimal number of devices for deploying a target model, as this relies on factors like model sizes and available hardware in different training scenarios. To address this, we introduce the concept of Model Placement Ratio, which is a value ranging from 0 to 1 that represents the proportion of devices allocated for a specific model. This provides enhanced adaptability. For instance, the Flattening strategy for a Reward model has a ratio of 1, indicating that the model is deployed on all devices. On the other hand, the Interleaving strategy for the Reward model, as depicted in Figure 4, has a ratio of 0.5, signifying that the model is deployed on half of the devices. Then the engine constructs a communication subgroup on the deployed devices of the Reward model for synchronization. It also belongs to a parent communication group that is connected with all models’ subgroups for data interaction. Then our Execution Engine generates the model placement mapping, which places the model to physical devices properly according to the model placement ratio, considering the device capacity and topology, such as intra- and inter-node. This mapping involves detailed ranking assignments for each model. Overall, this abstraction effectively shields users from the complexities associated with model placement, parallelism techniques, device computing capacity, and topology.
@@
Data Interaction. Data interaction caused by the placement strategy is also encapsulated within the Execution Engine. The Interleaving and Separation strategies introduce communication among different communication subgroups during the different training stages, and the APP handles these data interaction logics. Each training or prediction operation of the models is treated as an individual Operator. The rank assigned to each model performs an AllGather operation to collect the upstream data before executing the Operator. Once the Operator is executed, the resulting data is scattered back downstream.
@@
This abstraction enables a general interface and backend for creating a wide range of RLHF algorithms for research or industrial applications. Users can focus on the essence of RLHF modeling without being burdened by the intricacies of the underlying implementations.
V-C Adaptive Parallel Planner
Figure 6: The Flowchart of Adaptive Parallel Planner
The objective of the Adaptive Parallel Planner is to search for hyperparameters, including optimal model placement mapping, to maximize the training throughput of RLHF in various training scenarios while considering computing resource and model size constraints. Once the hyperparameters are determined according to the maximum overall floating-point operations per second (TFLOPS), the APP Execution Engines utilize it to execute the RLHF pipeline finally.
The workflow of the Adaptive Parallel Planner is illustrated in Figure 6. The input of the planner consists of NodeDesc (specifying the quantity, type, and memory of GPU devices), RLHFModelDesc (indicating the size of each model), and an optional Model Placement strategy which could be represented by Model Placement Ratio. The output includes hyperparameters, with the key model placement ratio being a crucial component. The hyperparameters also include batch size during the generation and training stages, as well as the Parallelism method (like ZeRO 1 or 2).
The workflow can be summarized as follows:
1. The MinResourceChecker initially calculates the minimum GPU resources required based on the NodeDesc, RLHFModelDesc, and optional Model Placement Ratio. If the available computing resources do not meet the minimum requirement, the process is halted.
2. Subsequently, the Planner explores the best configurations by using two search strategies. The first is a prior-based strategy that searches for the optimal configuration according to predefined rules. This strategy considers our prior knowledge, such as allocating resources ranging from 30% to 50% for Pred Actor models and training models respectively, while reserving the remaining resources for other prediction models. The second strategy is a grid search-based approach that exhaustively explores different combinations.
3. The Planner begins by iterating through a set of candidate model placement ratios for all four models. Each candidate placement ratio is represented as a vector, with each element corresponding to the model placement ratio. Next, the Planner proceeds to iterate through candidate hyperparameters for each model involved in the RLHF pipeline, including both the training and inference stages. In this step, NVIDIA NeMo[24] can be utilized to search for hyperparameters specific to each model. Once the hyperparameters for the RLHF pipeline are determined, they are passed on to the APP Execution Engine. This engine runs the training for a specific duration and calculates the TFLOPS.
4. Lastly, the APP engine outputs the resulting memory utilization and TFLOPS measurements to the Judger. It determines whether to continue the hyperparameter search based on predefined conditions, such as the specified iteration or duration of searching. The hyperparameters that yield the highest TFLOPS are then utilized in the production RLHF training process.
VI Experiments
@@
We conduct comparative experiments on both AC-NonShare and AC-Share model structures against the Flattening strategy used by DeepSpeed-Chat and trlX respectively. We evaluate the model throughput while using the maximum batch size available for each framework before Out-of-Memory (OOM) for fairness. These experiments encompass variations in model sizes, device scales, and device types, allowing us to assess the effectiveness of our APP framework across different training scenarios.
VI-A Experiment Settings
@@
LLM Backbones. We employ Llama[33] of different parameter sizes as our LLM backbone, including 7B, 13B, 33B, and 65B. For demonstration purposes, we keep consistent model size for each model involved in RLHF pipeline, but our APP framework could adapt to training scenarios where each RLHF model may have varying model sizes and other LLM models, such as Opt[42] and GLM[9].
@@
Baseline. We conduct a comparison between the APP framework and the current state-of-the-art methods in RLHF training: DeepSpeed-Chat with Hybrid Engine for the AC-NonShare case and trlX for the AC-Share case.
@@
VI-B Experiments and Analysis
We compare the throughput of our APP framework against DeepSpeed-Chat and trlX respectively. Specifically, in the training of the 7B model, we utilize ZeRO-2 for training models and ZeRO-0 for inference models. For the training of 13B, 33B, and 65B models, we employ ZeRO-3 for both stages. Furthermore, our Separation strategy incorporates tensor parallelism.
VI-B1 Comparing with DeepSpeed-Chat
Interleaving vs Flattening Strategy. We implement two Interleaving strategies to compare them with the Flattening strategy used in DeepSpeed-Chat. In $\text{Interleaving}_{1}$, the Reference and Reward models are placed on half of the GPU devices. In $\text{Interleaving}_{2}$, we interleave not only the Ref and Reward models but also the Actor and Critic models. Based on our experiments as illustrated in Table I, we conclude that the $\text{Interleaving}_{1}$ strategy is 12% faster than the Flattening strategy in DeepSpeed-Chat in small training scenarios, such as with 7B and 13B models. Additionally, the $\text{Interleaving}_{2}$ strategy is 71% faster than DeepSpeed-Chat in larger training scenarios, specifically with 65B models when employing ZeRO-3. The speedup ratio further increases when scaling out to more devices.
i) $\text{Interleaving}_{1}$ vs. Flattening Strategy. During the training of 7B size models using 1$\times$8 devices, we observe that $\text{Interleaving}_{1}$ strategy in our APP framework leads to a throughput increase from 16.12 to 17.35 samples per second, representing a speedup of 7.63% compared with Flattening strategy in DeepSpeed-Chat. This speedup ratio further improves to 12% when scaling out from 1$\times$8 to 4$\times$8 devices. The reason behind this improvement is that the adoption of the interleaving strategy effectively halved the memory redundancy of the Ref and Reward models. Consequently, the freed-up GPU memory can be utilized to expand the overall batch size, as explained in Section V-A. In this case, the Interleaving strategy achieves a maximum batch size of 36, whereas DeepSpeed-Chat’s maximum batch size is 24. However, it is important to note that $\text{Interleaving}_{1}$ does not perform well with larger model scales, such as in the 33B and 65B training scenarios. This is because these models require ZeRO-3 to reduce the memory redundancy to enable training large language models, while there is no memory redundancy of model states in this case.

### 25 Jan 2024 v2

@@ IV Our Framework
Figure 3: The timeline of (a) Flattening strategy vs. (b) Interleaving strategy. Two homogeneous GPU devices are utilized for demonstration. Under the Flattening strategy, both the Ref model and Reward model are deployed on both devices, whereas under the Interleaving strategy, the Ref model and Reward model are allocated to separate devices. As in the figure, the efficiency improvement of the interleaving strategy is attributed to two main reasons: i) Reduced memory redundancy: In the Flattening strategy (a), $4$ models including Actor, Critic, Ref and Reward are allocated on both GPU-$1$ and GPU-$2$, while in Interleaving strategy (b), the Ref model and Reward model are placed on GPU-$1$ and GPU-$2$ exclusively. This reduces the memory redundancy using low-level ZeRO Parallelism by reducing participating devices for the Reward model or Ref model from $2$ to $1$ as discussed in Section II-A. Then it allows us to increase the batch size to enhance memory utilization (as the figure illustrates, the batch size can be raised from $6$ to $12$). ii) Reduced communication cost: As the Interleaving strategy assigns the Ref model and Reward model to two separate GPU devices, it enables independent and parallel forward computation without communication between two devices in high-level ZeRO Parallelism.
IV-A Model Placement Strategies
IV-A1 Interleaving Strategy
@@
The main idea of the Interleaving strategy is to partially parallelize the pipeline by allocating the independent models on exclusive group of GPU devices, as well as shrinking the number of computing devices required for each model. We observe that either Reward&Ref Inference (and Actor&Critic training models) could operate independently, thereby it is possible to eliminate part of the serial execution. For example, the Ref model and Reward model both take the generated responses from the Actor model as input and output the intermediate results into the experience buffer without synchronization. Also, training Actor and Critic nodes have no dependency since they do not need synchronization during the training phase. This design helps the parallelization of independent models in RLHF training. This lowers the memory redundancy and shrinks the scale of participating devices used by independent models rather than using all devices in the Flattening strategy.
@@
We compare the distributed architecture and computation procedure between the Flattening strategy and our Interleaving strategy. We use a small number of devices as illustrative examples, while our strategy could be easily extended to large-scale computing devices. In terms of Model Placement, the Flattening strategy treats the four models as a unit and deploys them on all devices, as depicted in Figure  2(a). By contrast, our Interleaving strategy distinctly assigns the Ref model and Reward model to two distinct device groups (GPU 0-3 and 4-7) [^2] Our interleaving strategy could further place the Actor and Critic model on exclusive devices. We only illustrate the interleaving between Reward and Ref for simplicity., as illustrated in Figure  2(b). It reduces the number of devices by half used by either the Ref model or the Reward model. It’s important to note that responses generated from the Actor are located on GPUs where the Ref model is not placed since the Actor model is deployed on all GPU devices, as shown in the timeline in Figure  3(b). Therefore, it is necessary to perform two extra communication stages: an AllGather primitive to gather the generated results from all devices before forwarding, and then perform a Scatter primitive to distribute the results back to all devices. However, these two additional communication stages can be considered negligible by overlapping with computation. In other words, computation and communication occupy different cuda streams, allowing computation to begin upon receiving the first batch while the remaining batches can be communicated synchronously with the ongoing computation. We conclude that the Interleaving strategy shortens the time-cost generation duration from two aspects.
Figure 4: The timeline of Separation strategy on homogeneous (a) and heterogeneous GPU devices (b). As illustrated in the figure, we set up shadow critic and shadow actor models and placed them on separate GPU devices (as shown in figure (a) placed on GPU-3 and GPU-4, or in figure (b) placed on the inference-dedicated GPU-4 and GPUs 6 and 7), thereby decoupling the training and inference runtime in the pipeline. Additionally, in the heterogeneous Separation strategy in (b), these two stages are designated to specialized devices for training or inference purposes. As discussed in Section IV-A2, our Separation strategy could benefit from targeted optimization for each stage. Also, this strategy does not require waiting for the entire generation process to complete before proceeding with the Forward and Forward & Backward operation with pipeline parallelism shown in (b).
@@
Firstly, in terms of memory consumption, our Interleaving strategy reduces the number of participating devices for either the Ref or Reward model, effectively minimizing memory redundancy when applying ZeRO Parallelism. This allows our method to leverage the saved memory to increase the total batch sizes, resulting in higher overall throughput. For instance, when employing the Flattening strategy as displayed in Figure  3(a) under ZeRO-0, all devices hold replicas of the model states, e.g. of Reward Model, while fewer devices need to hold the replicas under our Interleaving strategy as depicted in Figure  3(b).
@@
Secondly, from a communication efficiency perspective, having fewer participating devices leads to reduced communication overhead in both the generation and training stages. It’s important to note that while increased levels of ZeRO parallelism reduce memory redundancy in model states, it comes at the cost of increased communication as discussed in SectionII-B. In this case, the Interleaving strategy can further leverage high-speed intra-node bandwidth and reduce inter-node communication by colocating the Reward model or Ref model on a single machine with usually 8 GPU devices, particularly in small to medium training scenarios, e.g. 7B or 13B models.
IV-A2 Separation Strategy
@@
To shorten the generation duration, our Separation strategy involves separating the training and the inference runtime of the RLHF pipeline. Specifically, we add extra shadow actor and critic models for inference only without a mixture usage of runtime. These shadow models are dedicated to inference and their parameters are synchronized with the corresponding training models periodically. This division not only allows us to run the training and generation models simultaneously but also leverages optimization techniques tailored for training or inference. It also could easily utilize the extra heterogeneous devices in clusters, since the high-end homogeneous GPU devices are scarce. Consequently, this approach significantly accelerates the generation stage and re-balance the workloads between these stages in the RLHF.
@@
This strategy significantly speeds up the entire RLHF training, though it introduces certain memory redundancy due to additional models and communication overhead. Figure  2(c) illustrates the placement of the training Actor and Critic models and inference models, including the Shadow Actor and Critic models, on separate devices. Optionally, we could further place the training Actor and Critic models on exclusive devices in an interleaving manner to further decrease the memory redundancy and communication costs. However, there is a need for communication for intermediate results and parameter synchronization between the training Actor/Critic models and the prediction models. Firstly, as shown in Figure  4(a), extra point-to-point communications are required before or after the Forward stage of each prediction model. Additionally, the training Actor and Critic models need to pull the intermediate results from the inference devices. These extra communication stages introduce some idle time per device, known as bubble overhead. However, these bubbles can be minimized by overlapping computation and communication stages through the use of pipeline parallelism as discussed in Section II-A. As illustrated in Figure  4(a), the bubbles are nearly hidden through micro-batching in a pipeline parallelism manner, especially with a higher number of generation times. Secondly, it is necessary to synchronize the trained model parameters between the training Actor/Critic models and shadow models after each round of PPO training. The synchronization overhead is inevitable but acceptable [^3] It is critical to clarify that this process is not asynchronous training, but still synchronous training without compromising model accuracy. It thanks to the off-policy property of RLHF that the training phase involves sampling from a buffer rather than tightly following the forward stage..
Although this approach introduces some complexity and overhead, it provides significant benefits from three different perspectives: 1) Our Separation strategy decouples training and inference runtime of the pipeline, enabling targeted optimization techniques for each stage, as discussed in Section III-A. For example, the time-consuming Actor Generation can be accelerated by utilizing specialized parallelism strategies that are well-suited for inference workloads, such as intra-node tensor parallelism by reducing low-bandwidth inter-node communication as discussed in II. 2) The strategy also facilitates the integration of evolving inference optimizations like quantization[34], which are less compatible with model training. There will be considerable runtime swapping overheads when using a mixture of training and inference runtime. 3) The strategy allows scaling out additional heterogeneous devices tailored for training or inference workloads, especially for the time-consuming generation stage. This flexibility in either quantity or type of devices offers significant adaptability to diverse training scenarios.
IV-B APP Execution Engine
The APP Execution Engine is designed to provide a user-friendly interface for modeling experts to easily build their own RLHF training pipeline while hiding the messy details of distributed computing and intermediate data interaction as illustrated in Figure 5. The APP Execution Engine encapsulates each model involved in the RLHF pipeline considering both AC-Share and AC-NonShare structures. Our modular design allows for the encapsulation of each model and corresponding distributed techniques. These techniques include the model placement strategy, singular model parallelism (such as ZeRO in DeepSpeed and Tensor Parallelism in Megatron), as well as various optimization techniques in other libraries. We further the abstraction in detail:
@@ IV-B APP Execution Engine
Model Placement Abstraction. While our placement strategies offer general guidelines on how to distribute these models, there remains a gap in determining the optimal number of devices for deploying a target model, as this relies on factors like model sizes and available hardware in different training scenarios. To address this, we introduce the concept of Model Placement Ratio, which is a value ranging from 0 to 1 that represents the proportion of devices allocated for a specific model. This provides enhanced adaptability. For instance, the Flattening strategy for a Reward model has a ratio of 1, indicating that the model is deployed on all devices. On the other hand, the Interleaving strategy for the Reward model, as depicted in Figure 3, has a ratio of 0.5, signifying that the model is deployed on half of the devices. Take the Reward Model in the model placement ratio representation shown in Figure 5, e.g. [1, 1, 0.5, 0.5], as an example. The engine constructs a communication subgroup on the deployed devices of the Reward model for synchronization. It also belongs to a parent communication group that is connected with all models’ subgroups for data interaction. Then our Execution Engine generates the model placement mapping, which places the model to physical devices properly according to the model placement ratio, considering the device capacity and topology, such as intra- and inter-node. This mapping involves detailed ranking assignments for each model.
@@
Hiding Intermediate Data Interaction. Data interaction caused by the placement strategy is also encapsulated within the Execution Engine. The Interleaving and Separation strategies introduce communication among different communication subgroups during the different stages as discussed in Section IV-A, and the APP engine handles these data interaction logics. Each training or forward operation of the models is treated as an individual Operator. The rank assigned to each model performs an AllGather operation to collect the upstream data before executing the Operator. Once the Operator is executed, the resulting data is scattered back downstream.
@@
Overall, this abstraction enables a general interface and backend for creating a wide range of RLHF algorithms for research or industrial applications. Users can focus on the essence of RLHF modeling without being burdened by the intricacies of the underlying details, as shown in Figure 5.
Figure 5: The User interface of APP Execution Engine.
V Experiments
@@
We conduct comparative experiments on both AC-NonShare and AC-Share model structures against the Flattening strategy used by DeepSpeed-Chat and trlX respectively. We evaluate the model throughput while using the maximum batch size available for each framework before Out-of-Memory (OOM) for fairness. These experiments encompass variations in model sizes, device scales, and device types, allowing us to assess the effectiveness and adaptability of our APP framework across different training scenarios.
V-A Experiment Settings
@@
LLM Backbones. We employ Llama[26] of different parameter sizes as our LLM backbone, including 7B, 13B, 33B, and 65B. For demonstration purposes, we keep consistent model size for each model involved in RLHF pipeline, but our APP framework could adapt to training scenarios where each RLHF model may have varying model sizes and other LLM models, such as Opt[35].
@@
Baseline. We conduct a comparison between the APP framework and the current state-of-the-art methods in RLHF training: DeepSpeed-Chat[33] with Hybrid Engine for the AC-NonShare case and trlX[3] for the AC-Share case.
@@
V-B Experiments and Analysis
We compare the throughput of our APP framework against DeepSpeed-Chat and trlX respectively. Specifically, in the training of the 7B model, we utilize ZeRO-2 for training models and DDP for inference models. For the training of more large models, e.g., 13B, 33B, and 65B models, we have to employ ZeRO-3 for both stages. Furthermore, our Separation strategy incorporates tensor parallelism.
V-B1 Comparing with DeepSpeed-Chat
Interleaving vs Flattening Strategy. We implement two Interleaving strategies to compare them with the Flattening strategy used in DeepSpeed-Chat. In $\text{Interleaving}_{1}$, the Ref and Reward models are placed on half of the GPU devices. In $\text{Interleaving}_{2}$, we interleave not only the Ref and Reward models but also the Actor and Critic models. Based on our experiments as illustrated in Table I, we conclude that the $\text{Interleaving}_{1}$ strategy is 12% faster than the Flattening strategy in DeepSpeed-Chat in small training scenarios, such as 7B model. Additionally, the $\text{Interleaving}_{2}$ strategy is 71% faster than DeepSpeed-Chat in larger training scenarios, specifically with 65B models when employing ZeRO-3. The speedup ratio further increases when scaling out to more devices.
i) $\text{Interleaving}_{1}$ vs. Flattening Strategy. During the training of 7B size models using 1$\times$8 devices, we observe that $\text{Interleaving}_{1}$ strategy in our APP framework leads to a throughput increase from 16.12 to 17.35 samples per second, representing a speedup of 7.63% compared with Flattening strategy in DeepSpeed-Chat. This speedup ratio further improves to 12% when scaling out from 1$\times$8 to 4$\times$8 devices. The reason behind this improvement is that the adoption of the interleaving strategy effectively halved the memory redundancy of the Ref and Reward models. Consequently, the freed-up GPU memory can be utilized to expand the overall batch size, as explained in Section IV-A. In this case, the Interleaving strategy achieves a maximum batch size of 36, whereas DeepSpeed-Chat’s maximum batch size is 24. However, it is important to note that $\text{Interleaving}_{1}$ does not perform well with larger model scales, such as in the 33B and 65B training scenarios. This is because these models require ZeRO-3 to reduce the memory redundancy to enable training large language models, while there is no memory redundancy of model states in this case.